OpenStudy (anonymous):

How do you find slope intercept form given two points? Ex: (28,40) and (30,0)

OpenStudy (jlastino):

You should find the slope first (y2-y1)/(x2-x1) x1= 28 x2=30 y1=40 y2=0 Then we will assign that value to be m now use the equation (y-y1)= m(x-x1) sub values into that equation and equate to "y" only and you should get y=mx+b

OpenStudy (anonymous):

I'm not very... quick with math so I understood the fire sentence. could you make the rest as easy as you can?

OpenStudy (anonymous):

first*

OpenStudy (jlastino):

Ok so we find the slope first (y2-y1)/(x2-x1) i assigned values for x1 x2 y1 and y2 x1= 28 x2=30 y1=40 y2=0 (0-40)/(30-28) -40/2 We can simplify this fraction by dividing -40 by 2 =-2 -2 will be our slope. the variable used for slope is the letter m m=-2 Now we will use another equation y-y1= m(x-x1) Remember the values that i assigned earlier y1=40 x1=28 m=-2 I substituted these values into the equation y-40=-2(x-28) to simplify this equation, i first distribute (meaning to multiply) -2 to (x-28) so it becomes -2x+56 (56 becomes positive because negative*negative = positive our equation so far: y-40=-2x+28 but we should make it look like the slope intercept form y=mx+b but what is b? we should equate our equation to y only(add 40 to both sides) y-40+40=-2x+28+40 y=-2x+68 and that's the slope intercept form of the equation :D where the slope is m=-2 and the y-intercept "b" is 68 :D

OpenStudy (anonymous):

That's so easy now! Thank you!